# How Much Can We Forget about Data Contamination?

### v3

@@ A.1 Additional Discussion of Data Contamination Assumptions and Setting
In this paper, we consider only exact contamination. This means we contaminate the training data exactly with the text the model is later evaluated on. In the literature, it has been shown that non-exact contamination (re-worded questions, translation into a different language) can affect benchmark performance, too. For example, Yang et al. [71] have shown that a 13B parameter Llama 2 Model [68] can achieve an accuracy increase of over 20 percentage points after training on re-phrased benchmark questions. We decided against considering non-exact contamination for this paper because the models we train from scratch are much smaller than those for which non-exact contamination results have been shown. This means these models are less capable of making sense of related information, potentially leading us to underestimate the effect of non-exact contamination for realistic training runs.
In addition, we consider contamination with individual benchmark questions, inserted into the training data at random positions. We consider this setup because we are interested in contamination from the perspective of leakage, where individual benchmark questions may enter the training data via different documents (for example, as quotes in Wikipedia articles, a case described in Brown et al. [10]). This contrasts with the setup where a dataset is present in the training data as a long contiguous string, which we conjecture might have a similar impact but be easier detectable [57]. The fact that we contaminate with benchmark questions also sets us apart from related works that study data contamination and memorization for random strings and uniquely identified objects [12, 13]. It is worth highlighting that the results between these two setups might differ, especially considering the time it takes to forget an example.
@@
A.2 Additional Details on Evaluation and How Contamination was Performed
Benchmark Questions and Evaluation. We use code from OLMo [30] to format the different benchmark questions. This code is again based in part on the EleutherAI Evaluation Harness [24]. The benchmark questions are multiple-choice, and the different options are presented to the model zero-shot as possible sentence continuations. The prediction is the sentence continuation with the largest likelihood. For the small GPT-3 models, we normalize by the number of tokens [23]. For OLMo, we rely on the evaluation framework that is part of the code repository.
Inserting benchmark questions into the training data. A batch of LLM training data consists of $B$ sequences of $S$ tokens, resulting in a batch size of $B\times S$. For example, OLMo-1B is trained with $B=2048$ and $S=2048$; the batch for a single gradient step contains $\sim$4M tokens [30]. Individual sequences in a batch usually contain multiple texts separated by a special end-of-text token. We insert benchmark questions at random positions into the pre-training data, separated at the beginning and end with the end-of-text token.
A.3 Filtering Near-Duplicate Benchmark Questions
Upon close inspection of the different benchmarks, it turns out that there are exact and near-duplicate questions both within and across benchmarks. Consider, for example, the following two examples from HellaSwag [72]:
and
Note that these are the ground-truth options of two different benchmark questions. Similar patterns can be observed for many benchmarks, for example, because a single text document was used to create different benchmark questions. Here is another example from PiQA [7]:
PiQA 1: Goal: how do you close a cabinet? Solution: push the door shut.
and
PiQA 2: Goal: how do you close a cabinet? Solution: shut the door.
@@ A.3 Filtering Near-Duplicate Benchmark Questions
(b) WinoGrande
(c) ARC-Easy
(d) PiQA
(a) Forgetting on Novel Data
(b) After 3 Epochs
(c) After 7 Epochs
Figure 8: Forgetting without weight decay. We perform the forgetting experiment from Section 4.2 without weight decay (instead of the default weight decay of 0.1). This figure depicts the same quantities as Figure 2 in the main paper. We see that there is significant forgetting without weight decay. At the same time, forgetting without weight decay occurs somewhat slower, especially for 144 times repeated contamination. In particular, the accuracy gaps depicted in (b) are slightly larger than the corresponding accuracy gaps in Figure 2, and the difference in (c) is statistically significant (compare Figure 2).
A.4 Reproducibility
@@
The results in this paper were obtained using the OLMo codebase, available at https://github.com/allenai/OLMo, and the llm.c codebase, available at https://github.com/karpathy/llm.c. Our code is fully reproducible, including the random positions at which benchmark questions were inserted into the training data. We trained on the 100BT split of the FineWeb-Edu dataset, available at https://huggingface.co/datasets/HuggingFaceFW/fineweb-edu. Code for this paper is available at https://github.com/tml-tuebingen/forgetting-contamination/. Model training relied on Pytorch [60] and was performed on 8xA100 nodes for all experiments except the continual pre-training of OLMo-7B that ran for a total of 3 weeks on an 8xH100 node.
A.5 Proof of Proposition 1

### v4

@@ A.3 Additional Discussion of Data Contamination Assumptions and Setting
In this paper, we consider only exact contamination. This means we contaminate the training data exactly with the text the model is later evaluated on. In the literature, it has been shown that non-exact contamination (re-worded questions, translation into a different language) can affect benchmark performance, too. For example, Yang et al. (2023) have shown that a 13B parameter Llama 2 Model (Touvron et al., 2023) can achieve an accuracy increase of over 20 percentage points after training on re-phrased benchmark questions. We decided against considering non-exact contamination in this paper because the models we train from scratch are much smaller than those for which non-exact contamination results have been shown. This means these models are less capable of making sense of related information, potentially leading us to underestimate the effect of non-exact contamination for realistic training runs.
In addition, we consider contamination with individual benchmark questions, inserted into the training data at random positions. We consider this setup because we are interested in contamination from the perspective of leakage, where individual benchmark questions may enter the training data via different documents (for example, as quotes in Wikipedia articles, a case described in Brown et al. (2020)). This contrasts with the setup where a dataset is present in the training data as a long contiguous string, which we conjecture might have a similar impact but be easier detectable (Oren et al., 2024). The fact that we contaminate with benchmark questions also sets us apart from related works that study data contamination and memorization for random strings and uniquely identified objects (Carlini et al., 2019, 2021). It is worth highlighting that the results between these two setups might differ, especially considering the time it takes to forget an example.
@@
A.4 Additional Details on Evaluation and How Contamination was Performed
@@
Benchmark Questions and Evaluation. We use code from OLMo (Groeneveld et al., 2024) to format the different benchmark questions. This code is again based in part on the EleutherAI Evaluation Harness (Gao et al., 2024). The benchmark questions are multiple-choice, and the different options are presented to the model zero-shot as possible sentence continuations. The prediction is the sentence continuation with the largest likelihood. For the small GPT-3 models, we normalize by the number of tokens (Gao, 2021). For OLMo, we rely on the evaluation framework that is part of the code repository.
Inserting benchmark questions into the training data. A batch of LLM training data consists of $B$ sequences of $S$ tokens, resulting in a batch size of $B\times S$. For example, OLMo-1B is trained with $B=2048$ and $S=2048$; the batch for a single gradient step contains $\sim$4M tokens (Groeneveld et al., 2024). Individual sequences in a batch usually contain multiple texts separated by a special end-of-text token. We insert benchmark questions at random positions into the pre-training data, separated at the beginning and end with the end-of-text token.
A.5 Filtering Near-Duplicate Benchmark Questions
Upon close inspection of the different benchmarks, it turns out that there are exact and near-duplicate questions both within and across benchmarks. Consider, for example, the following two examples from HellaSwag (Zellers et al., 2019):
and
Note that these are the ground-truth options of two different benchmark questions. Similar patterns can be observed for many benchmarks, for example, because a single text document was used to create different benchmark questions. Here is another example from PiQA (Bisk et al., 2020):
PiQA 1: Goal: how do you close a cabinet? Solution: push the door shut.
and
PiQA 2: Goal: how do you close a cabinet? Solution: shut the door.
A.6 Measuring the Overlap Between the Benchmark Questions and the Pre-Training Data
Here, we analyze the overlap between the 44000 de-duplicated benchmark questions and the pre-training data. Note that we did not de-contaminate the pre-training data (we use FineWeb-Edu “as-is”). In addition, note that our experimental design is valid even if the pre-training data is already contaminated (if there is already contamination, our experiments measure the causal effect of inserting the benchmark questions $k$-additional times). Our implementation of fuzzy string matching used to de-duplicate the benchmark questions is not efficient enough to match all benchmark questions against the pre-training data. Hence, we follow Singh et al. (2024) and report the length of the longest common substring between a benchmark question and the pre-training data. For simplicity, we report the overlap between all benchmark questions and the entire 100B token split of FineWeb-Edu. In our experiments, the training data is a random subset of these 100B tokens.
@@ A.6 Measuring the Overlap Between the Benchmark Questions and the Pre-Training Data
(b) PiQA
(c) Social-i-QA
(d) BoolQ
(e) MMLU
(f) WinoGrande
(g) ARC-Easy
Figure 7: Overlap between the benchmark questions and the 100B token split of FineWeb-Edu (Single Best Match). For every benchmark question, we search in the pre-training data for the longest common substring. The figure depicts the distribution of the length of the longest common substring, normalized by the length of the benchmark question, for the different benchmarks.
Figure 7 depicts the overlap between the benchmark questions and the pre-training data. For every benchmark question, we compute the length of the longest common substring divided by the length of the benchmark question (that is, a value of 1 means that the benchmark question occurs verbatim in the pre-training data). Figure 7 depicts the result of this computation in aggregated form. From Figure 7, we see that there is little overlap between the pre-training data and Hellaswag, PiQA, Social-i-QA, and Winogrande. For ARC-Easy, MMLU, and especially BoolQ, we find that there exist benchmark questions that have very high overlap with the pre-training data. Depending on the format of the questions and the nature of the overlap, some of these cases essentially correspond to verbatim contamination. In other cases, there might be a high overlap between the benchmark question and the pre-training data, but it is not necessarily clear to what degree the text in the pre-training data actually reveals the true answer to the benchmark question. On BoolQ, for example, many of the overlaps with the pre-training data occur for the descriptive contexts of the questions, which are the same for all answer options. Or consider the following example from ARC-Easy, where the pre-training data contains all four answer options.
Arc-Easy: Question: In Colonial America, people used ice to help keep foods fresh. They cut the ice from lakes and ponds during the winter and stored the ice in ice houses. They sometimes used hay as an insulator to prevent the ice from melting. If you wanted to build an icehouse today, which of the following would be the best material to use as an insulator?
Answer: foam blocks
and
Pre-training data: In Colonial America, people used ice to help keep foods fresh. They cut the ice from lakes and ponds during the winter and stored the ice in ice houses. They sometimes used hay as an insulator to prevent the ice from melting. If you wanted to build an icehouse today, which of the following would be the best material to use as an insulator?
(A) dried leaves
(B) foam blocks
(C) plastic wrap
(D) rock salt
In order to build a wooden table, […]
For the correct interpretation of Figure 7, note that is depicts the length of the single best match between the benchmark questions and the entire pre-training corpus of 100B tokens. For comparison, Figure 8 depicts the distribution of the length of the 10 best matches in the pre-training data. From Figure 8 (a), we see that searching for the 10 best matches does not affect the overall frequency distribution on Social-i-QA, for which there was no significant overlap in the first place. In contrast, for BoolQ, MMLU, and ARC-Easy, for which there was significant overlap, we observe that the frequency distribution of the 10 best matches shifts significantly to the left, indicating that regions of high overlap occur less than 10 times in the overall pre-training data.
In summary, we find that there are benchmark questions that have high overlap with the pre-training data, and that these questions are mostly from BoolQ, MMLU and ARC-Easy. At the same time, most of our 44000 benchmark questions do not have any significant overlap with the pre-training data as measured by the longest common substring metric. For those questions that have a high overlap, such an overlap usually occurs less than 10 times in the 100BT split of FineWeb-Edu. More research is needed to investigate different types of overlap between benchmark questions and the pre-training data, and whether they lead to benchmark overfitting (see also Jiang et al. (2024), who investigate n-gram-based overlap).
(a) Social-i-QA
(b) BoolQ
(c) MMLU
(d) ARC-Easy
Figure 8: Overlap between the benchmark questions and the 100B token split of FineWeb-Edu (10 Best Matches). For every benchmark question, we search in the pre-training data for the 10 longest common substrings. The figure depicts the distribution of the length of the longest common substrings, normalized by the length of the benchmark question, for the different benchmarks.
A.7 Reproducibility
@@
The code for this paper is available at github.com/tml-tuebingen/forgetting-contamination. It relies on the OLMo codebase, available at github.com/allenai/OLMo, and the llm.c codebase, available at github.com/karpathy/llm.c. Our code is fully reproducible, including the random positions in which the benchmark questions were inserted into the training data. We trained on the 100BT split of the FineWeb-Edu dataset, available at huggingface.co/datasets/HuggingFaceFW/fineweb-edu. Model training relied on Pytorch (Paszke et al., 2019) and was performed on 8xA100 nodes for all experiments except the continual pre-training of OLMo-7B, which ran for 6 weeks on 4xH100.
A.8 Proof of Proposition 1
@@
(a) HellaSwag
(b) WinoGrande
(c) ARC-Easy
(d) PiQA
(a) HellaSwag
(b) WinoGrande
(c) ARC-Easy
(d) PiQA
(a) Forgetting on Novel Data
(b) After 3 Chinchilla
(c) After 7 Chinchilla
Figure 11: Forgetting without weight decay. We perform the forgetting experiment from Section 4.2 without weight decay (instead of the default weight decay of 0.1). This figure depicts the same quantities as Figure 2 in the main paper. We see that there is significant forgetting without weight decay. At the same time, forgetting without weight decay occurs somewhat slower, especially for 144 times repeated contamination. In particular, the accuracy differences depicted in (b) are slightly larger than the corresponding accuracy differences in Figure 2, and the difference in (c) is statistically significant (compare Figure 2).